Insert operations and grouping symbols to make the following equations true:
3 3 3 3 3=1
3 3 3 3 3=9
I always get 3 when I try to answer this problem.
3×3÷3+3-3=3; ((3×3)+(3-3))÷3=3; and many more that I've tried.

Answer by onlinepsa(22)

(3 + 3 x 3) ÷ 3 - 3 =1
(3 - 3 ÷ 3) x 3 + 3 =9

Thanks,
PRD
https://onlinepsa.wordpress.com